# High-fidelity universal gates in the 171Yb ground state nuclear spin qubit

### v1

@@ I Introduction
In this work, we demonstrate manipulation of quantum states in the ground-state nuclear spin of 171Yb atoms through individually controlled single-qubit gate operations and two-qubit operations based on sequential state-selective coherent excitation to a long-lived clock state and Rydberg state. We estimate a controlled-Z (CZ) fidelity of 99.72(3)(99.40(3)%) with (without) post-selection from two-qubit Clifford Randomized Benchmarking (CRB) experiments, and a single-qubit CRB fidelity of 99.963(2)%. Using a benchmark sequence that is insensitive to single-qubit phases (relevant to situations where such phases can be canceled between pairs of gates by using echo techniques as in [13, 9]) and anti-symmetric errors, we infer a CZ gate fidelity of 99.84(6)(99.56(5)%) with (without) post-selection. The combination of these spatially selective high-fidelity gates with the previously demonstrated continuous loading [1] and mid-circuit measurement [7] is expected to lead to new demonstrations of quantum error correction and complex circuits among many physical qubits, which will be explored in future work [26].
@@ IV.1 Clock Shelving
Figure 2: Clock benchmark. (a) Probability $P_{1}$ of measuring an atom in state |$\ket{1}$, after repeated application of SCP $X^{\mathrm{clk}}_{\pi}$ pulses on the IZ, for initial state |$\ket{1}$. Here we measure 99.860(7)shelving fraction per pulse, per atom. (b) Spin locking measurement for the $X^{\mathrm{clk}}_{\pi/2}-Y^{\mathrm{clk}}(T)-X^{\mathrm{clk}}_{-\pi/2}$ circuit, where $Y^{\mathrm{clk}}(T)$ is a pulse along the geometrical Y axis in the |$\ket{1}$ |$\ket{c}$ Bloch sphere with duration $T$ and Rabi rate $\Omega^{\mathrm{Y}}_{\mathrm{SL}}$. For each $\Omega^{\mathrm{Y}}_{\mathrm{SL}}$, we infer the laser frequency noise $S^{\mathrm{SL}}_{\nu}(f)$ from the temporal decay in coherence. We also show the interferometer laser frequency noise $S^{\mathrm{int}}_{\nu}(f)$ (black). Data below 4 kHz is heavily impacted by acoustic noise present in the interferometer (blue shaded region). (c) Pre-selected (red circles), post-selected on atoms remaining in the qubit subspace (blue squares), and survival (green triangles) for a clock-GERB experiment ($U=X^{\mathrm{clk}}_{\pi}X^{\mathrm{clk}}_{\pi}$) measured in the two atom basis. Solid lines are fits to $ap^{x}+b$, where $x$ is the number of GERB blocks, and $b=1/9$ for the pre-selected curve, $b=1/4$ for the post-selected in qubit subspace curve, and $b=0$ for the loss curve (see SM). Error bars represent $1\sigma$ confidence intervals. Inset representing U-GERB sequence as described in the main text.
@@
We measure the characteristic clock-GERB curve for $U=X^{\mathrm{clk}}_{\pi}X^{\mathrm{clk}}_{\pi}$ in Figure 2(c). We notice that this measurement is subjected to errors from the eight 1Q gates used. We characterize those in the case where $U$ is the identity operation and measure a contribution of 0.32(2)per atomic pair per GERB block. For this set, characteristic of a well-tuned system, we measure a pre-selected fidelity for the shelve-unshelve clock sequence (red circles) of 99.80(1)per pair, after removing the 1Q error. The post-selected fidelity of the clock shelve-unshelve sequence on atoms that remain on the qubit subspace at the end of the circuit is 99.94(2)(blue squares), which is mostly affected by decoherence between the optical and ground state qubits. The difference between the pre- and post-selected fidelities points to leakage (atoms that remained in |$\ket{c}$), and loss. By recycling atoms from the 3P0 state and imaging them afterwards, we can measure a pair survival probability (green triangles) of 99.979(4)after each shelve-unshelve pulse (see SM).
@@ IV.3 Two-Qubit Gate Benchmarking
Figure 4: Two-qubit Clifford RB (a) and CZ-GERB (b) benchmarking, showing the probability of measuring atoms in their initial state $P_{00}$ as a function of gate depth. Blue (red) data points represent probability of returning to the initial two qubit state with (without) post-selecting on atoms that are in either nuclear spin qubit at the end of the circuit. In green, pair survival is plotted. Solid curves are fits to the function $ap^{x}+b$, where $x$ is the circuit depth, and $b=1/4$ for the post-selected in qubit subspace curve, and $b=0$ for the pre-selected and survival curves (see SM). Error bars represent $1\sigma$ confidence intervals.
@@
The excess pair loss (0.11(3)%) measured in CZ-GERB (Figure 4(b)) relative to clock-GERB (Figure 2(c)) contributes half of the additional pre-selected infidelity measured between the two metrics, 0.20(2)and 0.44(5)respectively. Errors in entangling or single-qubit phases between consecutive CZ gates will lead to excess infidelity. Table 1 summarizes the observed infidelities. We provide a more detailed final error budget supported by in-situ measurements and simulations in the SM, as well as details regarding the depth scans and fidelity extraction.
@@
Table 1: Infidelity contribution and pair loss per CZ gate as measured by GERB and CRB. Clock-GERB, CZ-GERB, and 2Q CRB infidelities are corrected by the error of a single 1Q gate and the average number of 1Q gates in each circuit.
@@ Appendix A System Details
Key aspects of our experimental system have been previously described in Refs. [7, 1]. In this work, all state preparation and measurement is performed in optical tweezers, not a cavity-enhanced optical lattice as in [1].
The lasers required to create both science and reservoir arrays, to rearrange atoms between reservoir trap sites, and for the Raman beams needed for the single-qubit rotations are combined on a dichroic mirror stack before being delivered to the vacuum chamber via a high numerical aperture microscope objective (NA = 0.65, field of view = 0.5 mm). Another similar objective is placed at the other side of the vacuum chamber, and is used exclusively to collect 556 nm scattered light in order to perform low-loss, state-selective imaging of individual atoms.
Both the reservoir and science traps are generated via SLM phase patterns [58] that are imaged on the microscope objective. The arrays used for this work consist of 84 trapping sites, distributed in 12 rows and 7 columns. The top row of each array, which we call the interaction zone (IZ), is displaced from the rest of the array by 6 $\mu$m. The distance between sites in the IZ row is 3 $\mu$m. The reservoir traps have radial trap frequencies (along the $x$ and $y$ directions) of $\omega_{\mathrm{res}}/(2\pi)=110$ kHz, while the science trap frequency is $\omega_{\mathrm{sci}}/(2\pi)=50$ kHz. Both arrays are spatially matched using camera measurements on a lower numerical-aperture monitoring system, and by realizing atomic transfer experiments to calibrate any systematic offsets. To transfer atoms from one array to another, we linearly ramp up the power of one potential as we ramp down the other over 2 ms. All gates used in this work are performed in the science array IZ, while imaging, cooling, and state preparation are performed in sites of the reservoir array at the same locations.
@@
[…]
To deterministically load atoms into the IZ, we rearrange atoms with a single 483 nm tweezer created by a pair of crossed acousto-optical modulators (AODs). For this work, we perform CZ gates on either one or two pairs of sites in the IZ. On instances where we load two atomic pairs, these pairs are separated by 12 $\mu$m. After the atoms are loaded, we image, cool and prepare them in state |$\ket{0}$ using global addressing. Once the atoms have been initialized, we transfer them to the 460 nm science traps. At this point, we run arbitrary circuits, i.e. GERB, Clifford RB, RPE calibrations, etc, as described in more detail in the following appendices.
In a typical gate characterization sequence in the IZ we use several readout images with intermediate ground state optical pumping and clock state repumping to determine if each atoms was in state |$\ket{0}$, |$\ket{1}$, |$\ket{c}$ or lost. Data is analyzed in the two-qubit computational basis. We then use some of these images to properly post-select the circuit success on some conditions. In this way we can analyze errors arriving from leakage, loss or anything that affects the qubit subspace.
@@ Appendix B Single-Qubit Rotations
We perform arbitrary local single-qubit rotations using orthogonally polarized Raman beams whose frequency and position are controlled by two pairs of crossed AODs [10]. Both Raman beams are combined in a polarizing beam splitter and delivered to the vacuum chamber through the high numerical-aperture objective. At the atom plane, the beams are focused to 1.2 $\mu$m $1/e^{2}$ radius. One beam is polarized along the magnetic field and the other orthogonal to it. The Raman pair is detuned by $\Delta_{1Q}/(2\pi)=-5$ GHz from the ${}^{3}\text{P}_{1},F=1/2$ manifold.
@@
In order to realize the correct parallel control, we utilize upstream acousto-optical (AOM) and electro-optical (EOM) modulators. Calibration of the $X_{\pi/2}$ gate requires calibration of the bare qubit frequency, the differential light-shift on the qubit states due to the addressing laser, and the pulse area, as well as careful alignment of the Raman beams to the atoms. We calibrate the bare qubit frequency (near 388.9 kHz) through a 100 ms Ramsey spectroscopy experiment in the science traps, where we typically observe changes of a 1 Hz scale daily. For this small array, we do not observe significant inhomogeneities in the measured frequency across the IZ sites.
@@
To calibrate the differential light-shifts and pulse area, we control the relative and common optical powers between the two Raman beams at a single site level. Both calibrations rely on repeated application of pulses, with or without phase changes between them, similarly as done in some Robust Phase Estimation (RPE) experiments. Raman beam alignments are realized by mapping individual beams AC Stark shift into phase changes in a Ramsey sequence, similarly as in [9]. Faster camera-based measurements between the trap arrays and the 1Q array are used to correct drifts from its set-points. Our single-qubit gate performance is not limited by fundamental processes such as intermediate state scattering. Rather, based on RPE measurements and camera-based measurements, we attribute a large contribution to this error to quasi-static drifts in the relative position between traps and 1Q beams. Simulations suggest that 100 nm drifts and misalignment can explain typically observed errors.
@@
When we implement any GERB sequence for a 2Q gate we rely on the 1Q gates to randomize the gate of interest over a sub-set of arbitrary states. The random Haar distributed rotations are constructed from a library of angles $\phi_{0},\phi_{1},\phi_{2}$, such that $R_{\mathrm{rand}}=Z[\phi_{0}]X_{\pi/2}Z[\phi_{1}]X_{\pi/2}Z[\phi_{2}]$ samples each qubit’s Hilbert space uniformly. At the end of the GERB sequence, we apply a single pre-computed rotation $R_{\mathrm{f}}=Z[\phi^{f}_{0}]X_{\pi/2}Z[\phi^{f}_{1}]X_{\pi/2}Z[\phi^{f}_{2}]$ to return atoms to the initial qubit state, provided $U$ is perfect. We benchmark our GERB sequence against the 1Q Clifford RB result, by running it with $U=Id$. Each GERB block consist of four $X_{\pi/2}$ gates per atom, so eight per pair. We typically measure 0.32(2)infidelity per GERB block (computed on an atomic pair for consistency with 2Q benchmarks), which is consistent with eight times the average gate infidelity measured in the Clifford RB sequence. To remove the 1Q error from other GERB measurements, i.e. when $U$ is not an idle operation, we remove 0.16(1)per $U$ infidelity on each atomic pair.
@@ Appendix C Clock Operations
The clock beam propagates along the $x$-direction and at the atom plane is elliptically shaped with waists (wy,wz) = (400 $\mu$m, 35 $\mu$m). The large magnetic field and polarization selectivity, see Figure 1(a), suppresses other excitation paths, or even differential qubit phase shifts. We operate our clock with Rabi rates $\Omega_{\mathrm{clk}}/(2\pi)$ between 3 kHz and 15 kHz, placing us in the resolved-sideband limit with respect to our radial science trap frequencies of $\omega_{\mathrm{sci}}/(2\pi)\approx 50$ kHz. In this regime, we are sensitive to both the effects of atomic motion and laser frequency noise, which are the main contributors to our shelving error. To mitigate frequency noise seen by the atoms, we stabilize the phase of the delivered clock light via a fiber noise cancellation setup that references the phase of the delivered light to the vacuum chamber  [59]. Optical power delivered to the atoms is also actively stabilized.
@@
We characterize the temperature of atoms participating in the gates using spectroscopy of the motional clock sidebands along $x$, a representative example of which is shown in Figure 6(a). In these conditions, a single (non-composite) Blackman-shaped pulse with a peak Rabi rate of $2\pi\times 7$ kHz shows shelving fidelities slightly above 99%, measured by repeated application of the clock pulse as in Figure 2(a). We have verified via RPE measurements that residual infidelity is not due to a coherent errors, but rather the Debye-Waller effect from trapped atoms at finite temperature [42]. Using pulse-area-robust composite pulses improves this shelving fraction, at the price of applying longer pulses with greater sensitivity to both quasi-static and time-varying laser frequency noise.
@@
The lifetime of ground state atoms in the science or reservoir traps is mostly limited by vacuum and intensity noise on the traps, and is at least 5 s. However, for atoms in the clock state, Raman scattering of the 460 nm light drives population out the 3P0 manifold. For science traps with $50$ kHz trap frequency, we have measured $T_{1}=1.06(5)$ s, see Figure 6(b). For the $X^{\mathrm{clk}}_{\pi}$ SCPs pulses used in the main text (260 $\mu$s duration), we estimate that this scattering accounts for an $0.019$% error per CZ gate averaged over possible pair qubits states (see error budget appendix). This still poses a challenge to achieve high-fidelity circuit operation due to idle loss in a metastable qubit architecture at this trap wavelength and depth [35, 15].
@@
To measure laser frequency noise at 578 nm, we use a scheme similar to the one described in [45]. We employ a 1 km SM optical fiber interferometer surrounded by sound-absorbing material. One of the interferometer arms is frequency shifted by a 200 MHz AOM. Any laser frequency noise contributes to the interferometer phase noise. We measure it by observing the phase noise around the 200 MHz carrier with a spectrum analyzer (SignalHound SM200B). A typical measurement is shown in Figure 6(c). At low frequencies, the laser noise is overwhelmed by acoustic noise present in the room (blue shaded region in Figure 6(c)), while we attribute the majority of the noise above 4 kHz to laser frequency noise.
@@
Atomic spin locking measurements, as in Figure 2(b), are our best estimation of frequency noise above 1 kHz. A typical spin locking experiments performs a Ramsey sequence that prepares and atom along the Y axis of the optical qubit (|$\ket{1}\leftrightarrow\ket{c}$) Bloch sphere, but applies a drive with duration $T$ and Rabi rate $\Omega^{\mathrm{Y}}_{\mathrm{SL}}$ along the Y axis [23, 38]. In the presence of laser frequency noise, the atomic coherence is displaced from the Y axis, such that a final $\pi/2$ pulse along the X axis can not return the population to either of the poles, as shown in Figure 6(d). From this data, we determine the $1/e$ decay rate $\Gamma_{\mathrm{SL}}$, taking into account contributions from $T_{1}$. In the linear response approximation, the decay rate $\Gamma_{\mathrm{SL}}$ is $\Gamma_{\mathrm{SL}}=2\pi^{2}S_{\nu}(\Omega^{\mathrm{Y}}_{\mathrm{SL}}/(2\pi))$, where $S_{\nu}(f)$ is the double-sided laser frequency noise power spectral density [23]. We repeat this measurement for different Rabi rates $\Omega^{\mathrm{Y}}_{\mathrm{SL}}$ to reconstruct $S_{\nu}(f)$.
@@
To improve data-rate for calibrations involving our clock laser, we perform many gates in series on the same atoms. In order to recycle population from the clock state, we use a repumper on the fast ${}^{3}\text{P}_{0}\leftrightarrow^{3}\text{D}_{1},F=\frac{1}{2}$ 1388 nm transition. This beam propagates along the $x$-direction and is linearly polarized. About 2% of the atoms that undergo a repumping cycle decay to the 3P2, which is either weakly or not trapped in our tweezers. In order to recycle atoms with low loss, we optically pump whatever remaining atoms are in the ground state to |$\ket{0}$, apply a high Rabi rate $YXY^{\mathrm{clk}}$ $\pi$-pulse on the |$\ket{1}-\ket{c}$ transition, and repump with a 1388 nm resonant pulse whatever fraction was left in |$\ket{c}$ after the circuit. This reduces loss significantly and helps speed up relevant calibrations.

### v2

@@ I Introduction
In this work, we demonstrate manipulation of quantum states in the ground-state nuclear spin of 171Yb atoms through individually controlled single-qubit gate operations and two-qubit operations based on sequential state-selective coherent excitation to a long-lived clock state and Rydberg state. We estimate a controlled-Z (CZ) fidelity of 99.72(3)(99.40(3)%) with (without) post-selection from two-qubit Clifford Randomized Benchmarking (CRB) experiments, and a single-qubit CRB fidelity of 99.963(2)%. Using a benchmark sequence that is insensitive to single-qubit phases (relevant to situations where such phases can be canceled between pairs of gates by using echo techniques as in [13, 9]) and anti-symmetric errors, we infer a CZ gate fidelity of 99.84(6)(99.56(5)%) with (without) post-selection. The combination of these spatially selective high-fidelity gates with the previously demonstrated continuous loading [1] and mid-circuit measurement [7] is expected to lead to new demonstrations of quantum error correction and complex circuits among many physical qubits, as we have recently shown in [26].
@@ IV.1 Clock Shelving
Figure 2: Clock benchmark. (a) Probability $P_{1}$ of measuring an atom in state |$\ket{1}$, after repeated application of SCP $X^{\mathrm{clk}}_{\pi}$ pulses on the IZ, for initial state |$\ket{1}$. Here we measure 99.860(7)shelving fraction per pulse, per atom. (b) Spin locking measurement for the $X^{\mathrm{clk}}_{\pi/2}-Y^{\mathrm{clk}}(T)-X^{\mathrm{clk}}_{-\pi/2}$ circuit, where $Y^{\mathrm{clk}}(T)$ is a pulse along the geometrical Y axis in the |$\ket{1}$ |$\ket{c}$ Bloch sphere with duration $T$ and Rabi rate $\Omega^{\mathrm{Y}}_{\mathrm{SL}}$. For each $\Omega^{\mathrm{Y}}_{\mathrm{SL}}$, we infer the laser frequency noise $S^{\mathrm{SL}}_{\nu}(f)$ from the temporal decay in coherence. We also show the interferometer laser frequency noise $S^{\mathrm{int}}_{\nu}(f)$ (black). Data below 4 kHz is heavily impacted by acoustic noise present in the interferometer (blue shaded region). (c) Pre-selected (red circles), post-selected on atoms remaining in the qubit subspace (blue squares), pair leakage (brown diamonds) and pair survival (green triangles) for a clock-GERB experiment ($U=X^{\mathrm{clk}}_{\pi}X^{\mathrm{clk}}_{\pi}$) measured in the two atom basis. Solid lines are fits to $ap^{x}+b$, where $x$ is the number of GERB blocks, and $b=1/9$ for the pre-selected curve, $b=1/4$ for the post-selected in qubit subspace curve, and $b=0$ for the loss and leakage curves (see SM Appendix  F). Error bars represent $1\sigma$ confidence intervals. Inset representing U-GERB sequence as described in the main text.
@@
We measure the characteristic clock-GERB curve for $U=X^{\mathrm{clk}}_{\pi}X^{\mathrm{clk}}_{\pi}$ in Figure 2(c). We notice that this measurement is subjected to errors from the eight 1Q gates used. We characterize those in the case where $U$ is the identity operation and measure a contribution of 0.32(2)per atomic pair per GERB block. For this set, characteristic of a well-tuned system, we measure a pre-selected fidelity for the shelve-unshelve clock sequence (red circles) of 99.80(1)per pair, after removing the 1Q error. The post-selected fidelity of the clock shelve-unshelve sequence on atoms that remain on the qubit subspace at the end of the circuit is 99.94(2)(blue squares), which is mostly affected by decoherence between the optical and ground state qubits. The difference between the pre- and post-selected fidelities points to leakage (atoms that remained in |$\ket{c}$), and loss.
Leakage error represents population left in the clock state after each shelve-unshelve sequence ($U$). We are able to measure this contribution using an additional readout step preceded by an clock repumping state, that effectively brings all of the clock remaining clock population back to the ground state. By estimating the fraction of pairs that are present in this readout image, compared to the image that only reveal atoms present in the qubit subspace, we determine that clock leakage (brown diamonds) contributes a 0.13(2)infidelity in our clock-GERB sequence. Finally, using all of the readout steps we can determine the pair survival probability (green triangles) of 99.979(4)after each shelve-unshelve sequence (see SM Appendix  C).
@@ IV.3 Two-Qubit Gate Benchmarking
Figure 4: Two-qubit Clifford RB (a) and CZ-GERB (b) benchmarking, showing the probability of measuring atoms in their initial state $P_{00}$ as a function of gate depth. Blue squares (red circles) represent probability of returning to the initial two qubit state with (without) post-selecting on atoms that are in either nuclear spin qubit at the end of the circuit. Pair leakage (survival) measurements are shown with brown diamonds (green triangles). Solid curves are fits to the function $ap^{x}+b$, where $x$ is the circuit depth, and $b=1/4$ for the post-selected in qubit subspace curve, and $b=0$ for the pre-selected, leakage and survival curves (see SM). Error bars represent $1\sigma$ confidence intervals.
@@
The excess pair loss (0.11(3)%) measured in CZ-GERB (Figure 4(b)) relative to clock-GERB (Figure 2(c)) contributes half of the additional pre-selected infidelity measured between the two metrics, 0.20(2)and 0.44(5)respectively. Furthermore, clock state leakage is determined to be the same as the one measured in our clock-GERB data set, at 0.12(3)per gate, per pair for CZ-GERB data and 0.14(2)for the 2Q CRB set. Additionally, errors in entangling or single-qubit phases between consecutive CZ gates will lead to excess infidelity. Table 1 summarizes the observed infidelities. We provide a more detailed final error budget supported by in-situ measurements and simulations in the SM, as well as details regarding the depth scans and fidelity extraction.
@@
Table 1: Infidelity contribution, pair leakage and pair loss per CZ gate as measured by GERB and CRB. Clock-GERB, CZ-GERB, and 2Q CRB infidelities are corrected by the error of a single 1Q gate and the average number of 1Q gates in each circuit.
@@ Appendix A System Details
Key aspects of our experimental system have been previously described in Refs. [7, 1]. The lasers required to create both reservoir and science arrays, to rearrange atoms between reservoir trap sites, and the Raman beams needed for the single-qubit rotations are combined on a dichroic mirror stack. These beams are then delivered to the vacuum chamber via a high numerical aperture microscope objective (NA = 0.65, field of view = 0.5 mm). Another similar objective is placed at the other side of the vacuum chamber, and is used exclusively to collect 556 nm scattered light in order to perform low-loss, state-selective imaging of individual atoms. In this work, all state preparation and measurement operations are performed in optical tweezers, not a cavity-enhanced optical lattice as in [1].
The reservoir and science traps are each generated via SLM phase patterns [58] that are imaged on a microscope objective. The arrays used for this work consist of 84 trapping sites, distributed in 12 rows and 7 columns. The top row of each array, which we call the interaction zone (IZ), is displaced from the rest of the array by 6 $\mu$m. The distance between sites in the IZ row is 3 $\mu$m. The reservoir traps have radial trap frequencies (along the $x$ and $y$ directions) of $\omega_{\mathrm{res}}/(2\pi)=110$ kHz, while the science trap frequency is $\omega_{\mathrm{sci}}/(2\pi)=50$ kHz. Both arrays are spatially matched using camera measurements on a lower numerical-aperture monitoring system, and by realizing atomic transfer experiments to calibrate any systematic offsets. To transfer atoms from one array to another, we linearly ramp up the power of one potential as we ramp down the other over 2 ms. All gates used in this work are performed in the science array IZ, while imaging, cooling, and state preparation are performed in sites of the reservoir array at the same locations.
@@
After imaging, atoms are further cooled in the reservoir traps of the interaction zone to facilitate high-fidelity gate operations. In this work we use a variant of gray molasses cooling by addressing the atoms simultaneously with the $\sigma^{+}$ and $\pi$ polarized 556 nm beams addressing the ${}^{1}\textrm{S}_{0},m_{F}=\mp\frac{1}{2}\leftrightarrow^{3}\textrm{P}_{1},F=\frac{3}{2},m_{F}=\frac{1}{2}$ transitions. The $\sigma^{+}$ beam is generated by the same system as the imaging beam shown in Figure 1(a), while two $\pi$-polarized beams propagate along the $y$ and $z$ directions [1, 8]. These $\pi$-polarized beams are circular Gaussian beams with waists of 150 $\mu$m at the atom plane. The single-photon detuning of these beams is + 2.5 MHz from the excited state, while the frequency difference between the $\sigma^{+}$ and $\pi$ beams is equal to the qubit frequency of 389 kHz. After a cooling duration of 4 ms, we typically observe $\bar{n}=0.25(10)$, and do not observe significant heating or loss during the transfer of atoms between reservoir and science arrays.
To deterministically load atoms into the IZ, we rearrange atoms with a single 483 nm tweezer created by a pair of crossed acousto-optical modulators (AODs). For this work, we perform CZ gates on either one or two pairs of sites in the IZ. On instances where we load two atomic pairs, these pairs are separated by 12 $\mu$m. After the atoms are loaded, we image, cool and prepare them in state |$\ket{0}$ using global addressing. Once the atoms have been initialized, we transfer them to the 460 nm science traps. At this point, we run arbitrary circuits, e.g. GERB, Clifford RB, and RPE calibrations, as described in more detail in the following appendices.
In a typical gate characterization sequence in the IZ we use several readout images with intermediate ground state optical pumping and clock state repumping to determine if each atoms was in state |$\ket{0}$, |$\ket{1}$, |$\ket{c}$, or lost. Data is analyzed in the two-qubit computational basis. We then use some of these images to properly post-select the circuit success based on some conditions, or to identify population recovered by optical pumping and repumping. In this way we can analyze errors arriving from leakage, loss, or anything that affects the qubit subspace.
@@ Appendix B Single-Qubit Rotations
We perform arbitrary local single-qubit rotations using two orthogonally polarized Raman beams whose frequency and position are controlled by two pairs of crossed AODs [10]. The Raman beams are combined on a polarizing beam splitter and delivered to the vacuum chamber through a high numerical-aperture objective. At the atom plane, the beams are focused to 1.2 $\mu$m $1/e^{2}$ radius. One beam is polarized along the magnetic field and the other orthogonal to it. The Raman pair is detuned by $\Delta_{1Q}/(2\pi)=-5$ GHz from the ${}^{3}\text{P}_{1},F=1/2$ manifold.
@@
Upstream acousto-optical (AOM) and electro-optical (EOM) modulators enable parallel control of single-qubit operations. Calibration of the $X_{\pi/2}$ gate requires calibration of the bare qubit frequency, the differential light-shift on the qubit states due to the addressing lasers, and the pulse area, as well as careful alignment of the Raman beams to the atoms. We calibrate the bare qubit frequency (near 388.9 kHz) through a 100 ms Ramsey spectroscopy experiment in the science traps, where we typically observe changes of $\sim 1$ Hz daily. We do not observe significant inhomogeneities in the measured qubit frequency across the IZ sites.
@@
Calibration of the differential light-shifts and pulse area are performed by controlling the relative and common optical powers between the two Raman beams at a single site level. Both calibrations rely on repeated application of pulses, with or without phase changes between them, similar to RPE experiments. Raman beam alignments are realized by mapping an individual beam’s AC Stark shift into phase changes in a Ramsey sequence, similar to Ref. [9]. Faster camera-based measurements between 1Q array and the trap arrays are used to correct drifts from its setpoint. Our single-qubit gate performance is not limited by fundamental processes such as intermediate state scattering. Rather, based on RPE and camera-based measurements, we attribute the majority of the 1Q error to quasi-static drifts in the relative position between traps and 1Q beams. Simulations suggest that 100 nm drifts and misalignment can explain typically observed errors.
@@
GERB sequences for the 2Q gate rely on 1Q operations to randomize the gate of interest over a subset of arbitrary input states. Random Haar distributed rotations are constructed from random angles $\phi_{0},\phi_{1},\phi_{2}$ with distributions weighted such that $R_{\mathrm{rand}}=Z[\phi_{0}]X_{\pi/2}Z[\phi_{1}]X_{\pi/2}Z[\phi_{2}]$ samples each qubit’s Hilbert space uniformly. At the end of the GERB sequence, we apply a single pre-computed rotation $R_{\mathrm{f}}=Z[\phi^{f}_{0}]X_{\pi/2}Z[\phi^{f}_{1}]X_{\pi/2}Z[\phi^{f}_{2}]$ to return atoms to the initial qubit state, provided $U$ is perfect. We benchmark our GERB sequence against the 1Q Clifford RB result, by running it with $U=Id$. Each GERB block consist of four $X_{\pi/2}$ gates per atom, so eight per pair. We typically measure 0.32(2)infidelity per GERB block (computed on an atomic pair for consistency with 2Q benchmarks), which is consistent with eight times the average gate infidelity measured in the Clifford RB sequence. To remove the 1Q error from other GERB measurements, i.e. when $U$ is not an idle operation, we subtract 0.16(1)per $U$ infidelity on each atomic pair.
@@ Appendix C Clock Operations
The clock beam propagates along the $x$-direction and is elliptically shaped at the atom plane with waists (wy,wz) = (400 $\mu$m, 35 $\mu$m). The large magnetic field and polarization selectivity, see Figure 1(a), suppress other excitation paths, and also suppress differential qubit phase shifts. We operate our clock with Rabi rates $\Omega_{\mathrm{clk}}/(2\pi)$ between 3 kHz and 15 kHz, placing us in the resolved-sideband limit with respect to the radial science trap frequencies of $\omega_{\mathrm{sci}}/(2\pi)\approx 50$ kHz. In this regime, we are sensitive to both the effects of atomic motion and laser frequency noise, which are the main contributors to the shelving error. To mitigate frequency noise seen by the atoms, we stabilize the phase of the delivered clock light via a fiber noise cancellation setup that references the phase of the delivered light to the vacuum chamber [59]. Optical power delivered to the atoms is also actively stabilized.
@@
We characterize the atom temperature using spectroscopy of the motional clock sidebands along $x$, as shown in Figure 6(a). At typical temperatures, a single (non-composite) Blackman-shaped pulse with a peak Rabi rate of $2\pi\times 7$ kHz shows shelving fidelities slightly above 99%, measured by repeated application of the clock pulse as in Figure 2(a). We have verified via RPE measurements that residual infidelity is not due to a coherent errors, but rather the Debye-Waller effect from trapped atoms at finite temperature [42]. Using pulse-area-robust composite pulses improves this shelving fraction, at the price of applying longer pulses with greater sensitivity to both quasi-static and time-varying laser frequency noise.
@@
The lifetime of ground state atoms in the science or reservoir traps is mostly limited by vacuum and intensity noise on the traps, and is at least 5 s. However, for atoms in the clock state, Raman scattering of the 460 nm light drives population out the 3P0 manifold. For science traps with $50$ kHz trap frequency, we have measured $T_{1}=1.06(5)$ s, see Figure 6(b). For the $X^{\mathrm{clk}}_{\pi}$ SCPs pulses used in the main text (260 $\mu$s duration), we estimate that this scattering accounts for an $0.019$% error per CZ gate averaged over possible pair qubits states (see Appendix G). We note that such Raman scattering would act as an idle loss in metastable qubit architectures, posing a challenge to achieve high-fidelity circuit operation at this trap wavelength and depth [35, 15].
@@
To measure laser frequency noise at 578 nm, we use a scheme similar to the one described in [45]. We employ a 1 km SM optical fiber interferometer surrounded by sound-absorbing material. One of the interferometer arms is frequency shifted by a 200 MHz AOM. Laser frequency noise is mapped (via the transfer function of the delay line) into the interferometer beat note at 200 MHz, which is measured with a spectrum analyzer (SignalHound SM200B). A typical measurement is shown in Figure 6(c). At low frequencies, the laser noise is overwhelmed by acoustic noise present in the room (blue shaded region in Figure 6(c)), while we attribute the majority of the noise above 4 kHz to laser frequency noise.
@@
Atomic spin locking measurements, as in Figure 2(b), are our best estimation of frequency noise above 1 kHz. A typical spin locking experiments performs a Ramsey sequence that prepares an atom along the Y axis of the optical qubit (|$\ket{1}\leftrightarrow\ket{c}$) Bloch sphere, then applies a drive with duration $T$ and Rabi rate $\Omega^{\mathrm{Y}}_{\mathrm{SL}}$ along the Y axis [23, 38]. In the presence of laser frequency noise, the atomic coherence is displaced from the Y axis, such that a final $\pi/2$ pulse along the X axis can not return the population to either of the poles, as shown in Figure 6(d). From this data, we determine the $1/e$ decay rate $\Gamma_{\mathrm{SL}}$, taking into account contributions from $T_{1}$. In the linear response approximation, the decay rate $\Gamma_{\mathrm{SL}}$ is $\Gamma_{\mathrm{SL}}=2\pi^{2}S_{\nu}(\Omega^{\mathrm{Y}}_{\mathrm{SL}}/(2\pi))$, where $S_{\nu}(f)$ is the double-sided laser frequency noise power spectral density [23]. We repeat this measurement for different Rabi rates $\Omega^{\mathrm{Y}}_{\mathrm{SL}}$ to reconstruct $S_{\nu}(f)$.
@@
To improve data-rate for calibrations involving our clock laser, we perform many experiments in series on the same atoms, which can lead to leakage into the clock state. In order to recycle population from the clock state, we use a repumper on the fast ${}^{3}\text{P}_{0}\leftrightarrow^{3}\text{D}_{1},F=\frac{1}{2}$ 1388 nm transition. This beam propagates along the $x$-direction and is linearly polarized. However, about 2% of the atoms that undergo a repumping cycle decay to the 3P2, which is either weakly or not trapped in our tweezers, leading to loss. Therefore, before repumping, we optically pump atoms are in the ground state to |$\ket{0}$, apply a high Rabi rate $YXY^{\mathrm{clk}}$ $\pi$-pulse on the |$\ket{1}-\ket{c}$ transition, finally apply a resonant 1388 nm pulse to the clock state. This reduces loss significantly, speeding up relevant calibrations.